# AI-Assisted Game Management Decisions: A Fuzzy Logic Approach to Real-Time Soccer Substitutions

### v3

@@ 3.1.1 Problem Formulation
The Predictive Ceiling of Machine Learning: Current computational approaches to substitutions rely heavily on Supervised Machine Learning (SML) models trained on historical data. As noted by Mohandas et al. mohandas, these models achieve a prediction accuracy plateau of approximately 70%. This ”ceiling” exists because SML models are designed to mimic human behavior, thereby learning and replicating the cognitive biases of coaches (e.g., status quo bias or delaying defensive changes). Consequently, these models validate historical decisions rather than optimizing future outcomes, failing to identify necessary substitutions that deviate from conservative human norms.
@@ 3.1.2 Research Hypotheses
  • H2 (Prescriptive Hypothesis): A Fuzzy Logic-based Decision Support System, acting as an objective auditor of performance variables (Performance, Fatigue, Discipline), can identify high-risk tactical scenarios (e.g., ”invisible” defensive liabilities) earlier and more accurately than human intuition or predictive models trained on historical behavior.
  • H3 (Contextual Hypothesis): The modulation of disciplinary risk by tactical role (e.g., weighting a yellow card more heavily for a defender than a forward) significantly alters the substitution priority score, aligning the algorithm’s output with expert tactical consensus in critical defensive scenarios.
@@ 3.3.4 Data Validation
Before integration into the fuzzy system, consistency checks were performed. The recalculation of playerank_acumulativo_media_raw in the final stage (Stage 3) ensures that all cumulative metrics are based on the robust expanding().mean() operator applied to the complete and harmonized temporal trace of each player. This guarantees that all cumulative scores represent valid in-game performance dynamics and that the fuzzy system operates on stable, correctly aggregated data.
@@ 3.5.4 Inference and Final Calculation
4 Results and Discussion
Figure 3: (each point represents the interval up to the next, e.g., 0 = 0-5 min, 5 = 5–10 min, 10 = 10–15 min)
To assess the ecological validity and practical utility of the Fuzzy Control System (FCS) in a high-stakes environment, the model was applied to historical data from the 2018 FIFA World Cup quarter-final match between Brazil and Belgium. This case study was selected due to the critical nature of the tactical decisions required under pressure (with Brazil trailing 0-2 at halftime) and the availability of extensive post-match expert analysis to serve as ground truth for validation.
4.1 Adherence to Technical Commission Decisions
The simulation demonstrated a significant correlation between the Substitution Priority ($P_{final}$) generated by the model and the actual alterations made by the coaching staff. This alignment suggests that the fuzzy inference engine successfully identified performance bottlenecks in real-time. Table 4 summarizes the critical decision windows, comparing the model’s output—specifically the ranking of the player within the team at that specific minute—with the real-world outcome.
Table 4: Comparison between Model Priority ($P_{final}$) and Real Tactical Decisions (Brazil vs. Belgium, 2018).
Player Analysis Slice($t$) $P_{\text{final}}$ in the slice Rank at $t$ Predominant Fuzzy Logic Real Decision
Willian 40’-45’ 72.0 4th Rapid Drop (R04) Substituted out
Gabriel Jesus 55’-60’ 99.1 2nd Ineffective Forward (R08) Substituted out
Paulinho 60’-65’ 93.1 2nd Ineffective Midfielder Substituted out
Fagner 45’-85’ 100.0 1st Fatigue Not Substituted
4.1.1 Analysis of Executed Substitutions
The model correctly flagged the three players substituted by the head coach, Tite, as high-priority candidates for removal, validating the specific fuzzy rules designed for performance decay:
  • Willian (45’): At halftime, the system assigned a priority of $P_{final}=72.0$. While he was ranked 4th in substitution urgency, the system triggered the Rapid Drop rule (R04), identifying a sharp decline in technical actions during the final 20 minutes of the first half. The coach’s decision to substitute him immediately supports the model’s diagnosis of offensive ineffectiveness, later described by media analysis as a performance ”below expected levels” ESPN_2018.
  • Gabriel Jesus (58’): Approaching the hour mark, the model calculated a critical priority of 99.1, placing him as the 2nd highest priority in the squad. The system activated rule R08 (Ineffective Forward), triggered by the combination of a low cumulative performance score ($P_{cum}$) and a null value for goals and assists. Post-match statistics revealed the player had only 12 ball touches in 58 minutes Epoca_2018, empirically validating the fuzzy system’s ability to detect tactical isolation.
  • Paulinho (73’): With a priority of 93.1 till the minute 65 (last on-field activity registered), the model identified ”Early Fatigue” combined with low structural participation. This aligns with external critiques noting the player ”did not make his presence felt” and failed to provide the necessary offensive surprise Gazeta_2018.
4.2 The Counterfactual Case: The Value of Decision Support
The most significant finding lies in the divergence between the model and human decision-making regarding the right-back, Fagner. This discrepancy highlights the potential value of the FCS as a bias-mitigating tool.
Throughout the critical window (minutes 40 to 85), Fagner consistently held the Maximum Priority (100.0), ranking as the #1 candidate for substitution in the team, surpassing all players who were actually removed. The fuzzy inference engine identified a ”Critical Risk” scenario composed of:

### v5

@@ 4.2.1 Problem Formulation
The Predictive Ceiling of Machine Learning: Current computational approaches to substitutions rely heavily on Supervised Machine Learning (SML) models trained on historical data. As noted by mohandas, these models achieve a prediction accuracy plateau of approximately 70%. This ”ceiling” exists because SML models are designed to mimic human behavior, thereby learning and replicating the cognitive biases of coaches (e.g., status quo bias or delaying defensive changes). Consequently, these models validate historical decisions rather than optimizing future outcomes, failing to identify necessary substitutions that deviate from conservative human norms.
@@ 4.2.2 Research Hypotheses
  • H2 (Prescriptive Auditing Hypothesis): A prescriptive decision support system, acting as an objective auditor of performance and contextual variables (e.g., performance trends, fatigue, disciplinary exposure), can identify high-risk tactical scenarios—such as latent defensive liabilities—earlier and more reliably than human intuition or data-driven predictive models trained on historical behavior.
  • H3 (Contextual Modulation Hypothesis): The modulation of disciplinary risk by tactical role (e.g., weighting a yellow card more heavily for a defender than a forward) significantly alters the substitution priority score, aligning the system’s output with expert tactical consensus in critical defensive scenarios.
  • H4 (Prescriptive AI Formalization Hypothesis): The empirical validation of a prescriptive, auditing-based decision support system in a high-stakes environment supports the formalization of a distinct Prescriptive AI paradigm, in which the system objective is not to predict human decisions, but to support normative, risk-aware judgment through explicit reasoning and accountability.
@@ 4.4.4 Data Validation
Before integration into the fuzzy system, consistency checks were performed. The recalculation of playerank_acumulativo_media_raw in the final stage (Stage 3) ensures that all cumulative metrics are based on the robust expanding().mean() operator applied to the complete and harmonized temporal trace of each player. This guarantees that all cumulative scores represent valid in-game performance dynamics and that the fuzzy system operates on stable, correctly aggregated data. The experimental design yields high statistical density: the inference engine executes approximately 360 distinct decision audits (2 teams × 10 outfield players × 18 temporal windows). This granular approach allows us to evaluate the system’s stability across diverse game states (winning/losing, high/low fatigue) within a controlled adversarial environment, ensuring internal validity beyond a simple match-outcome correlation
@@ 4.6.4 Inference and Final Calculation
4.7 A Normative Taxonomy of Action-Oriented AI Systems: Formalizing Prescriptive AI
The ability to recommend or evaluate actions has been addressed across multiple research traditions and industrial domains under diverse labels, including decision optimization, action recommendation, prescriptive analytics, and AI-driven decision support. These approaches are often discussed as if they constituted a coherent class of systems. In reality, they occupy a heterogeneous and only loosely structured conceptual space, within which crucial distinctions regarding uncertainty, accountability, and human agency are frequently overlooked.
This section establishes a coherent conceptual hierarchy for action-oriented decision systems. We first characterize the broad and heterogeneous superset of action-oriented systems. We then situate Prescriptive Analytics as a structured decision-oriented framework within that space. Finally, we formally delineate Prescriptive Artificial Intelligence (Prescriptive AI) as a normative, high-stakes subset of Prescriptive Analytics, defined not by algorithmic choices but by its epistemological role in human decision-making.
4.7.1 Action-Oriented Decision Systems: A Heterogeneous Superset
Action-oriented decision systems form a broad superset of computational approaches whose shared property is the capacity to recommend, prioritize, or execute actions. This superset includes, among others:
  • Optimization and mathematical programming systems;
  • Simulation-based and what-if analysis tools;
  • Rule-based and expert systems;
  • Reinforcement Learning (RL) agents;
  • Recommendation and next-best-action systems;
  • AI-augmented decision support tools embedding predictive models into workflows.
Despite their shared action-producing nature, these systems differ substantially in objectives, assumptions, and interaction with human decision-makers. Many are designed primarily for automation or outcome maximization, often assuming stable objectives, well-defined environments, and limited requirements for interpretability or accountability. Consequently, action-oriented systems do not constitute a single paradigm, but rather a conceptually heterogeneous collection of approaches.
4.7.2 Prescriptive Analytics as a Decision-Oriented Subset
Within this heterogeneous landscape, Prescriptive Analytics provides a more structured and widely recognized framework. Originating in business intelligence and operations research, prescriptive analytics focuses on determining what actions should be taken to achieve specified objectives, typically through optimization, simulation, and rule-based reasoning Davenport2007; Lepenioti2020; Bertsimas2020.
Prescriptive Analytics narrows the scope of action-oriented systems by explicitly addressing decision recommendation. However, it remains largely agnostic with respect to how uncertainty, human judgment, or accountability should be handled. It encompasses both automated and decision-support systems and does not inherently distinguish between optimizing outcomes and supporting human deliberation.
Accordingly, Prescriptive Analytics is best understood as a decision-oriented subset of action-oriented systems: analytically grounded, but normatively under-specified.
Action-Oriented AI Systems (The Heterogeneous Superset) Reinforcement Learning (Automation-Oriented) Recommender Systems Predictive Workflows Prescriptive Analytics (Decision-Oriented) Black-Box Optimization Simulation-Based Planning Prescriptive AIDecision AuditingIntrinsic InterpretabilityRisk-Aware Reasoning
Figure 5: A normative taxonomy of action-oriented AI systems. The diagram illustrates the hierarchical relationship where Action-Oriented Systems form the superset. Prescriptive Analytics sits within as a decision-focused subset, while Prescriptive AI serves as the normative core, defined by auditability and interpretability requirements.
4.7.3 Formal Definition of Prescriptive AI
Within Prescriptive Analytics, we define Prescriptive Artificial Intelligence (Prescriptive AI) as a specialized normative subset designed explicitly for high-stakes, human-in-the-loop decision-making.
Definition. Prescriptive AI is a class of intelligent systems whose primary function is to audit, justify, and support human decisions under uncertainty, rather than to automate actions or predict future outcomes.
Prescriptive AI systems are not characterized by the techniques they employ, but by the role they play in the decision process. In particular, they are distinguished by the following defining properties:
  • A primary function of decision auditing rather than decision automation;
  • Explicit reasoning under uncertainty and asymmetric risk;
  • Intrinsic interpretability as a structural requirement, not a post-hoc feature Rudin2019;
  • Contestable recommendations that preserve human agency;
  • Evaluation criteria that decouple state assessment from outcome realization, explicitly avoiding outcome bias BaronHershey1988.
In its lexical sense, a “prescription” is the action of laying down authoritative rules or directions, or more generally, something prescribed as a rule mw:prescription. By this definition, a prescriptive statement is inherently normative: it asserts what ought to be done rather than merely predicting or suggesting an outcome. Consequently, any system that cannot be inspected, audited, or contested cannot properly prescribe.
Taken together, these properties establish a necessary boundary: a system that cannot be audited cannot prescribe; it can only suggest or automate. These properties impose normative constraints that exclude many systems commonly labeled as prescriptive. Prescriptive AI therefore constitutes a proper subset of Prescriptive Analytics, defined by epistemological and accountability requirements rather than by optimization or learning paradigms.
4.7.4 Decision-Making as Epistemic State Transition
The normative necessity of Prescriptive AI follows from a dynamic conception of rationality. In the framework of logical dynamics developed by van Benthem, decision-making is not primarily evaluated by realized outcomes, but by the epistemic state transitions induced by actions under partial and evolving information (vanBenthem2007; vanBenthem2011). Rationality, under this view, is a property of how an agent updates beliefs, constraints, and commitments at the moment of action, rather than of the stochastic realization of downstream consequences.
This perspective provides a formal grounding for the core design principles of Prescriptive AI. If decision quality is defined at time $t$, prior to outcome realization at time $t+n$, then systems designed to support or evaluate human decisions must operate on the informational state available at the moment of commitment. Predictive accuracy and ex post outcome optimization are therefore insufficient as normative criteria. A prescriptive system must instead audit whether a decision constitutes a coherent, risk-aware, and justifiable epistemic update given the available evidence and constraints.
From this standpoint, Prescriptive AI can be understood as an operational instantiation of epistemic action auditing. Rather than attempting to predict or automate behavior, the system evaluates the internal consistency of the decision state itself, including the alignment between observed signals, inferred risk, and permissible actions. This aligns directly with the logical-dynamic view in which actions are epistemic interventions that transform an agent’s informational state.
While the present work does not implement the full formal machinery of Dynamic Epistemic Logic, it adopts this foundational stance by treating decision evaluation as a state-based rather than outcome-based problem. The logical dynamics literature does not prescribe a specific algorithmic realization; instead, it characterizes the object of evaluation. Prescriptive AI derives its epistemic effectiveness precisely from this characterization: by auditing epistemic transitions rather than stochastic outcomes, it enables principled assessment of decision quality under uncertainty, irreversibility, and asymmetric risk.
4.7.5 Prescriptive AI vs. Predictive and Outcome-Driven Systems
Predictive AI systems, typically trained via supervised learning, map contextual inputs to historical human decisions or observed outcomes. While effective for forecasting and pattern discovery, such models inherently learn and reproduce historical behavior, including embedded cognitive biases such as status quo bias, sunk cost fallacy, and outcome bias.
Prescriptive AI systems, by contrast, operate as decision auditors. Rather than imitating past decisions, they evaluate the current state against explicit objectives, constraints, and domain knowledge. Their purpose is to assess whether a decision is justified given the information available at time $t$, not whether it coincidentally leads to a favorable outcome at time $t+n$.
This distinction is critical in stochastic, high-stakes environments. Evaluating decision systems based on realized outcomes introduces outcome bias. A high-risk decision state is not retroactively validated by a favorable random outcome. Prescriptive AI explicitly decouples state evaluation from outcome realization, treating disagreement with historical decisions or subsequent outcomes not as error, but as potential identification of latent risk.
Dimension Predictive AI Prescriptive Analytics Prescriptive AI
Primary Objective Forecast events or actions Recommend actions to optimize objectives Support normative human decision-making
Learning Paradigm Supervised / statistical learning from historical data Optimization, simulation, rules, or learned models Rule-based, symbolic, or hybrid (learning-agnostic)
Optimization Target Accuracy / loss minimization Expected utility or objective maximization Risk mitigation / decision justification
Relation to Bias Learns and reproduces historical bias Implicitly reflects modeled objectives Audits, constrains, and exposes bias
Evaluation Criterion Predictive accuracy Outcome-based performance State-based validity / risk detection
Outcome Dependence Strong (labels and metrics) Strong or implicit Explicitly decoupled from outcomes
Interpretability Optional / post-hoc Often secondary Structural requirement
Role of Human Source of labels Executor or overseer Accountable decision-maker
Table 7: Conceptual comparison between predictive AI, prescriptive analytics, and prescriptive AI paradigms. Prescriptive AI is distinguished not by a specific learning paradigm, but by its normative role in auditing and supporting human decisions under uncertainty.
4.7.6 Boundary Cases Within Prescriptive Analytics
Several classes of systems fall within Prescriptive Analytics but do not satisfy the defining criteria of Prescriptive AI:
Optimization-based systems compute actions that maximize objective functions under constraints Bertsimas2020. While effective in structured settings, they typically assume stable objectives and do not audit human decision justification under uncertainty.
Reinforcement Learning systems prescribe actions by optimizing expected reward SuttonBarto2018. Despite their prescriptive nature, they are often opaque and automation-oriented, limiting contestability and interpretability, and are particularly vulnerable to outcome bias in stochastic environments.
Recommendation and next-best-action systems prioritize actions based on engagement or conversion metrics, relying on outcome-driven evaluation and rarely addressing accountability or asymmetric risk.
Rule-based and expert systems offer interpretability but often rely on static knowledge and lack mechanisms for uncertainty-aware temporal auditing.
These systems are prescriptive in an analytic sense, but do not fulfill the normative role required of Prescriptive AI.
4.8 Hybrid Systems and Overlapping Set Membership
The conceptual sets described here are not mutually exclusive. Hybridization is the norm in modern AI systems, which frequently combine symbolic reasoning, optimization, statistical learning, and deep learning.
As with the relationship between machine learning and deep learning, subset membership reflects conceptual scope rather than exclusivity. A system may employ optimization or learning techniques associated with Prescriptive Analytics while simultaneously satisfying the normative criteria that qualify it as Prescriptive AI. Classification therefore depends not on the techniques employed, but on the system’s functional role in the decision-making process.
4.8.1 Cross-Domain Applicability: From Pitch to Critical Operations
While elite-level soccer serves as an observable laboratory for adversarial decision-making, the structural constraints of the domain—namely irreversibility, severe time pressure, and asymmetric risk—are isomorphic to those encountered in several safety-critical sectors. The prescriptive auditing framework proposed herein passos2025 does not rely on domain-specific mechanics, but on the temporal auditing of decision-relevant performance signals and agent state evolution. Consequently, the architecture is transferable to environments in which human judgment is susceptible to fatigue, status quo bias, or desensitization under sustained cognitive load.
  • Clinical Triage and High-Dependency Care: In emergency and critical-care settings, clinicians operate under persistent cognitive load, where subtle signs of deterioration may be overlooked due to desensitization or alarm fatigue. A prescriptive auditor could operate over continuous indicators of physiological state evolution—analogous to the performance signals used in this study—to surface latent risk trajectories that remain within nominal ranges but exhibit adverse momentum. Rather than producing diagnostic predictions, the system would prescribe prioritized re-evaluation, supporting timely intervention before irreversible deterioration occurs.
  • Financial Risk Management (Trading Desks): In financial decision-making, human operators are known to exhibit systematic biases such as sunk cost effects and loss aversion, leading to prolonged exposure to unfavorable positions. A prescriptive auditing system could continuously evaluate the alignment between risk exposure and market dynamics, issuing normative recommendations when exposure exceeds logic-based thresholds. In this sense, the system functions not as a predictor of market outcomes, but as an auditor of decision states under volatility.
  • Aviation and Control Room Operations: In air traffic control and other safety-critical control rooms, operator fatigue and cognitive degradation constitute latent failure modes that are difficult to observe directly. A prescriptive auditor could assess indicators of cognitive performance degradation—such as response latency or interaction patterns—to recommend preemptive role rotation or task redistribution before safety margins are compromised. The “substitution” logic instantiated in this work thus generalizes to proactive management of human operational capacity.
Across these domains, the core function of the system remains invariant: to decouple the evaluation of the decision state at time $t$ from stochastic outcomes at time $t+n$, thereby mitigating outcome bias in high-variance, high-stakes environments.
4.8.2 Terminological Convergence and Alternative Taxonomies
Current developments in machine intelligence exhibit a structural pivot from purely predictive architectures toward systems designed to constrain, guide, or audit agentic behavior. This convergence reflects the growing recognition that stochastic outcome prediction is frequently insufficient for evaluating performance in high-stakes, dynamic environments where process validity and contextual coherence supersede result optimization. We introduce the term Prescriptive AI to describe this emergent class of systems. Unlike predictive models that estimate the probability of a future state, Prescriptive AI establishes a normative framework for evaluating the coherence of a decision at the precise moment of its execution.
This perspective aligns with foundational work on the logical dynamics of information and agency, particularly van Benthem’s account of decision-making as an epistemic state transition under uncertainty (vanBenthem2007; vanBenthem2011). While the present work does not implement the formal machinery of dynamic epistemic logic, it adopts this conceptual stance by evaluating decisions based on the informational state and normative criteria available at the moment of action, rather than on the stochastic realization of downstream outcomes. By treating decision-making as a dynamic process of information uptake and evaluation, Prescriptive AI analytically decouples the agent’s reasoning from stochastic environmental variability, enabling principled auditing of decision quality independently of realized results.
This definition allows us to position Prescriptive AI precisely relative to adjacent taxonomies. While we adopt this term to emphasize the normative function of the system, the framework intersects with, and provides empirical grounding for, several established conceptual traditions:
  • Decision Intelligence (DI): Commonly used in industrial and organizational contexts, Decision Intelligence emphasizes the engineering of decision workflows. The proposed framework constitutes a specific instantiation of DI in which the critical decision interface is mediated by intrinsically interpretable symbolic reasoning, addressing auditability requirements that are often under-specified in conventional DI implementations.
  • Normative AI: Within the ethics and governance literature, systems that encode explicit standards of acceptable behavior are often described as normative. The system employed in this work functions as a normative agent insofar as it enforces a codified standard of decision quality—derived from both hard performance constraints and soft contextual heuristics—against which human actions are evaluated in real time.
  • Cognitive Orthotics: In human–computer interaction, cognitive orthotics refer to systems designed to compensate for specific human cognitive limitations, such as attentional lapses or systematic biases. By explicitly correcting for effects such as status quo bias and outcome bias, the proposed framework operates as a real-time cognitive stabilization mechanism in high-stakes decision contexts.
  • Real-Time Algorithmic Auditing: Algorithmic auditing is traditionally framed as a post-hoc activity concerned with fairness or compliance. In contrast, the present approach extends auditing into the decision loop itself, enabling continuous, run-time evaluation of human decision states prior to the execution of irreversible actions.
Within this landscape, Prescriptive AI serves as a unifying descriptor that emphasizes the system’s functional output—the prescription—while retaining the requirement for intrinsic interpretability and auditability that is central across these literatures.
4.8.3 Implications for High-Stakes Decision Making
By explicitly positioning Prescriptive AI as a normative subset of Prescriptive Analytics, and Prescriptive Analytics as a structured subset of a broader heterogeneous space of action-oriented systems, this work resolves a persistent source of conceptual confusion in both academic and industrial discourse Sun2020; Wissuchek2022.
This hierarchy clarifies that:
  • Not all action-oriented systems are prescriptive analytics;
  • Not all prescriptive analytics systems are Prescriptive AI;
  • Hybrid systems may legitimately belong to multiple conceptual sets.
Prescriptive AI should therefore be understood not as a generic label for action recommendation, but as a distinct AI paradigm designed to audit and support human judgment under uncertainty. In the remainder of this work, elite soccer is employed not merely as a case study, but as a demanding stress-test environment to operationalize and validate these principles under time pressure, incomplete observability, and asymmetric risk.
4.9 Ethical Use of AI in High-Stakes Decision Systems
Auditing artificial intelligence systems in high-stakes decision-making contexts is not merely a technical challenge but a fundamentally ethical one. As argued by Rudin2019, deploying opaque, black-box models in domains where decisions have irreversible consequences is inherently problematic. When such systems are extended to environments with intense public scrutiny and continuous human judgment—such as elite professional soccer—the ethical risks are amplified rather than mitigated.
In contrast to low-stakes or background automation, in-game tactical decisions are subject to immediate visual auditing by multiple human observers, including coaches, players, analysts, referees, and millions of spectators. Any divergence between an algorithmic recommendation and human intuition is instantly questioned. In this setting, the adoption of black-box models poses a substantial risk of rejection, driven by two central ethical concerns: epistemic opacity and institutional distrust.
Epistemic opacity and public accountability.
The vast majority of stakeholders involved in professional sports—including coaches, athletes, and fans—do not possess formal training in artificial intelligence, let alone in complex models such as deep neural networks. When a black-box system produces a recommendation that contradicts human judgment, the absence of a transparent rationale forces stakeholders to accept the decision on faith alone. This requirement for blind trust is ethically untenable in high-stakes environments, where decisions can alter careers, financial outcomes, and competitive integrity.
In contrast, auditable and intrinsically interpretable models—such as fuzzy logic systems—enable post-hoc and real-time explanation of decisions at multiple levels of abstraction. In the proposed framework, the entire reasoning pipeline remains open to inspection: from the microscopic computation of performance scores, through fuzzy membership functions, to the macroscopic aggregation of linguistic rules that generate the final substitution priority. This transparency substantially reduces epistemic asymmetry between the system and its users, fostering informed trust rather than blind acceptance. Importantly, such auditability allows non-technical stakeholders to understand why a recommendation was made, even without understanding *how* the underlying mathematics operates.
Institutional distrust and risk of misuse.
High-stakes domains with significant financial flows are particularly vulnerable to suspicion, manipulation, and perceived conflicts of interest. Professional sports exemplify this risk. The rapid expansion of online betting markets has reshaped the economic landscape of elite soccer, with betting companies now acting as major sponsors of clubs and leagues. Concurrently, multiple high-profile cases of match manipulation and betting-related misconduct involving professional players have eroded public trust.
In such an environment, the deployment of opaque AI systems introduces severe ethical hazards. Consider a scenario in which a black-box model recommends substituting a star player during a decisive match, followed by a negative outcome. Without an explicit and intelligible justification, the decision becomes vulnerable to accusations of corruption, collusion with betting interests, or deliberate sabotage. Coaches would lack defensible grounds for their actions, players could be unfairly penalized without explanation, and clubs would face reputational and legal risks. These dynamics create strong institutional incentives to avoid black-box systems altogether, regardless of their predictive accuracy.
By contrast, prescriptive and interpretable decision support systems mitigate these risks by design. Rather than replacing human judgment, they aim to *augment* it, offering structured, explainable insights that the human decision-maker may accept, reject, or contextualize. This human-in-the-loop paradigm significantly reduces the probability of catastrophic errors while preserving accountability. Moreover, because every recommendation can be traced to explicit rules and contextual indicators, the system provides a defensible audit trail that protects coaches, players, and institutions alike.
Prescriptive AI as an ethical design principle.
The ethical challenge in high-stakes AI is not merely accuracy but responsibility. Predictive models trained to imitate historical decisions inevitably inherit human biases, including status quo bias, outcome bias, and risk aversion. In contrast, prescriptive systems grounded in symbolic reasoning evaluate current conditions against normative rules, enabling systematic auditing of human decisions rather than their replication.
This work adopts interpretability and prescriptiveness as core ethical design principles. By prioritizing transparency, auditability, and human agency, the proposed system aligns with emerging arguments that intrinsically interpretable models are not only preferable but necessary in domains where errors carry disproportionate consequences (rudin). Rather than automating authority, the system enhances human decision-making, offering a practical pathway for deploying AI responsibly in high-stakes environments.
Ultimately, while the adoption of AI in decision support is unavoidable, its ethical deployment is not optional. In contexts such as elite soccer—where decisions are public, consequential, and irreversible—only transparent, auditable, and human-centered systems can achieve legitimate and sustainable integration.
5 Acknowledgments
In the spirit of Open Science, the complete source code, including pre-processing pipelines, fuzzy inference engines, and validation scripts, is publicly available at https://github.com/Pedro-Passos77/AI-Assisted-Substitution-Decisions-A-Fuzzy-Logic-Approach-to-Real-Time-Game-Management. This manuscript constitutes a substantially expanded version of the research accepted by the Wharton Sports Analytics Journal and the accepted in the AAAI 2026 Bridge on LMreasoning) Passos2025_Wharton passos2026, with improvements focused on methodological robustness, theoretical foundation, proposed paradigm and extended case study. We gratefully acknowledge Luca Pappalardo and collaborators for making the Soccer match event dataset publicly available (Pappalardo2019), which was essential for this study. Generative AI tools were used to assist in the drafting and structural organization of this manuscript under strict human supervision; the authors remain fully responsible for the final content and its normative justification.
5.1 Limitations and Future Work
While the proposed Fuzzy Control System (FCS) demonstrates substantial utility as a prescriptive auditing tool, it is essential to delineate its epistemic scope and operational boundaries. The system is not designed to perform causal intervention, predict exogenous events, or replace human judgment. Its function is strictly limited to auditing the observable decision space given the information available at time $t$, and should therefore be interpreted as a normative support mechanism rather than an autonomous decision-maker.
From a data perspective, the reliance on minutes_played as a proxy for physical fatigue assumes a quasi-linear accumulation of workload, serving as an estimation rather than a direct physiological measurement. This abstraction does not capture inter-individual physiological differences (e.g., genetics), variations in match intensity such as the volume of high-speed running, or metabolic expenditure under sustained defensive pressure. As a result, the model may underestimate fatigue in high-exposure contexts or overestimate wear in structurally low-intensity tactical systems.
At the modeling level, the fuzzy rule base is statically defined using expert heuristics. While this design choice is central to intrinsic interpretability and contestability, it may limit generalization across divergent tactical philosophies (e.g., high-pressing versus possession-oriented systems) without manual recalibration of membership functions or rule weights ($\alpha$). This reflects a broader trade-off between explainability and adaptive flexibility that is intrinsic to symbolic reasoning systems.
Importantly, the framework does not account for abrupt exogenous events such as acute injuries, referee decisions, or random disruptions, which occur independently of observable fatigue or performance decay. Such events lie outside the epistemic reach of any decision-auditing system grounded in pre-intervention signals.
Future iterations of this framework aim to address these limitations through three complementary research directions: